Engineers must consider the breadths of male heads when designing helmets. The company researchers have determined that the population of potential clientele have head breadths that are normally distributed with a mean of 5.6-in and a standard deviation of 1-in. Due to financial constraints, the helmets will be designed to fit all men except those with head breadths that are in the smallest 0.5% or largest 0.5%.

What is the minimum head breadth that will fit the clientele? min =

What is the maximum head breadth that will fit the clientele? min
Answer #1

Solution:-

Given that,

mean = \mu = 5.6 in.

standard deviation = \sigma = 1 in.

Using standard normal table,

P(Z < z) = 0.5%

= P(Z < z ) = 0.005

= P(Z < -2.58 ) = 0.005  

z = -2.58

Using z-score formula,

x = z * \sigma + \mu

x = -2.58 * 1 + 5.6

x = 3.02 in.

min. 3.02 in.

Using standard normal table,

P(Z > z) = 0.5%

= 1 - P(Z < z) = 0.005  

= P(Z < z) = 1 - 0.005

= P(Z < z ) = 0.995

= P(Z < 2.58 ) = 0.995  

z = 2.58

Using z-score formula,

x = z * \sigma + \mu

x = 2.58 * 1 + 5.6

x = 8.18 in.

max. 8.18 in.